OpenStudy (anonymous):

Evaluate the indefinite integral of: x^4tanx dx I would like a hint please.

OpenStudy (anonymous):

Is the tan(x) a part of the exponent?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It is (x^4)(tanx) dx

OpenStudy (anonymous):

I think the indefinite integral can only be expressed as an infinite series.

OpenStudy (anonymous):

Is this a homework question or something?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Interesting. I feel that I've seen this question before, but my instinct now tells me that it can only be evaluated as a series. Let's see... y=(x^4)tan(x), ln(y)/ln(x)=4+ln(tan(x))... No, this approach doesn't work...